Assume you just purchased 100 shares of Apple stocks at $580.  You are worrying that the competition from other tablet PC producers will have a negative impact on Apple stock prices in 1 month. Generally speaking, you are still quite bullish on Apple stock.  

    (i)  In order to hedge against this downside risk, you establish a protective put position by buying a put option contract with around 1-month maturity on Apple stock (note: 1 contract consists of 100 options). Each put option has a strike price of $580 and a premium of $21.00. If AAPL’s stock price is $500 after 1 month, please compute the profit/gain from the overall protective put position (that consists of a stock and a put).

     (ii) If you feel the premium of the put option with strike price at $580 is quite expensive, what can we do to reduce the cost of protective put position?  Please explain the benefit/cost of this alternative way to limit the downside risk?  (It is an openquestion.)
Answer #1

i) If Apple's stock is at $500 after 1 month ,

profit from the stock = final value - initial value = 500-580 = - 80

profit from the put option = final value - initial value = ( 580-500) - 21 = 59

Oerall profit/gain = -80+59 = -$21 per option

i.e a total of -21*100 = -$2100 for 1 contract (100 options)

ii) As the option with strike price of $580 is costly, we can go for a lower strike price put option with a lower premium.This will help us lower the cost of protective put position (benefit). For example purchasing a put with a strike price of $550 can save us a lot of premium amount. This option may be available at around $10-$12 . However, in this case , the option becomes worthless if the prices rise above $550. Particularly, if price is in 550-580 range we lose money on both the stock as well as option.